# Number Cookbook: Number Understanding of Language Models and How to Improve It

### v1

@@ Abstract
Large language models (LLMs) can solve an increasing number of complex reasoning tasks while making surprising mistakes in basic numerical understanding and processing (such as $9.11>9.9$). The latter ability is essential for tackling complex arithmetic and mathematical problems and serves as a foundation for most reasoning tasks, but previous work paid little attention to it or only discussed several restricted tasks (like integer addition). In this paper, we comprehensively investigate the numerical understanding and processing ability (NUPA) of LLMs. Firstly, we introduce a benchmark covering four common numerical representations and 17 distinct numerical tasks in four major categories, resulting in 41 meaningful combinations in total. These tasks are derived from primary and secondary education curricula, encompassing nearly all everyday numerical understanding and processing scenarios, and the rules of these tasks are very simple and clear. Through the benchmark, we find that current LLMs fail frequently in many of the tasks. To study the problem, we train small models with existing and potential techniques for enhancing NUPA (such as special tokenizers, PEs, and number formats), comprehensively evaluating their effectiveness using our testbed. We also finetune practical-scale LLMs on our proposed NUPA tasks and find that 1) naive finetuning can improve NUPA a lot on many but not all tasks, and 2) surprisingly, techniques designed to enhance NUPA prove ineffective for finetuning pretrained models. We further explore the impact of chain-of-thought techniques on NUPA. Our work takes a preliminary step towards understanding and improving NUPA of LLMs. Our benchmark and code are released at https://github.com/GraphPKU/number_cookbook.
@@ 1 Introduction
The mathematical and reasoning abilities of large language models (LLMs) are currently quite impressive (OpenAI, 2023; Meta, 2024; OpenAI, 2024a; Yang et al., 2024a), capable of solving problems at the level of a graduate student or even more difficult ones like olympiad-level problems (He et al., 2024), GAOKAO (a nationwide examination of high school students applying to universities in China) (Zhang et al., 2024) and college mathematics (Tang et al., 2024). However, upon closer examination of the models’ outputs, we found that although the models demonstrate remarkable proficiency in problem-solving approaches, they often struggle with basic numerical understanding and processing — like a careless student who claims, “I know how to do it, but I didn’t get it right.” Some of these errors are quite surprising, such as believing that $9.11>9.9$ or making mistakes in simple addition $8/7+3/5$. These errors are a major cause of hallucinations when dealing with math, reasoning, and data analysis tasks, as the model presents seemingly correct problem-solving approaches, but ultimately produces incorrect results. Therefore, investigating and improving the fundamental “numerical understanding and processing abilities” (NUPA) of models is crucial.
@@
A more severe issue is that the numerical content is often deliberately simplified in these datasets. In various exam questions (like in the American Invitational Mathematics Examination (Li et al., 2024)), to focus on assessing students’ understanding of mathematical concepts — such as how to set up the correct equations and apply the right theorems — the numbers in both the questions and answers are often specially chosen to be integers. However, this is not the case in real-world scenarios (Chen et al., 2022a).
@@
On this benchmark, we rigorously test several state-of-the-art LLMs containing GPT-4o (OpenAI, 2024a), Llama-3.1 (Meta, 2024) and Qwen2 (Qwen Team, 2024). We ask models to directly output the answers without calling external tools. Although the latest LLMs perform well on some easiest tasks, their performance declines significantly as tasks become slightly more complex (such as multiplication, modulo operations, or digit-based calculations), or as the representation of numbers extends beyond basic integers. See Figure 2 of Section 2.4. The overall unsatisfactory performance highlights a pronounced mismatch between the claimed strong mathematical reasoning abilities and the poor practical, everyday numerical understanding and processing abilities of today’s LLMs.
@@
Finally, we discuss the potential of using chain-of-thought (CoT) techniques for numerical processing. Although CoT methods can break down complex problems into simpler sub-tasks and significantly increase the likelihood of obtaining correct answers, their drawbacks — such as consuming a large context window and requiring extended processing time — become particularly apparent in numerical tasks. We test a general CoT method known as RFFT (Hu et al., 2024), and find that for more complex tasks (such as those with $O(n^{2})$ complexity, including multiplication, division, and fraction simplification), chain-of-thought methods face scalability challenges, making them difficult to be applied in practical scenarios. It is noteworthy that in this paper, we do not discuss tool use methods (Schick et al., 2023; Lu et al., 2023) for NUPA as 1) we want to study the self-contained NUPA of LLMs, 2) calling external tools whenever encountering numbers increases the inference latency (Xu et al., 2024), and 3) we believe NUPA without tools is a necessary ability of AGI.
@@ 2 NUPA test: a benchmark for number understanding and processing ability
In this section, we will introduce our NUPA benchmark from the following four aspects: number representations, tasks, metrics, and result analysis of current LLMs. We will explain the rationale behind the inclusion (or exclusion) of specific representations and tasks in our benchmark, highlighting their distinctive features. By comparing these representations and tasks with previous benchmarks, we will demonstrate that several common tasks and abilities have been overlooked in prior evaluations.
@@ 2.1 Number representation
As we discuss above, we believe that the educational curricula on the primary and secondary school levels is a valuable reference for determining the essential NUPAs that LLMs should master. We identify four number formats that are both common and sufficient to cover most practical scenarios.
  • Integer: Integer is the first set of numbers that we met in primary school and the most common number representation. Some experiments on animals (Davis & Memmott, 1982) and infants (Strauss & Curtis, 1981) have found that understanding of (simple) integers is somewhat innate because it is used naturally in counting various objects. Understanding and operations of more complex integers are also acquired very early in human civilization (Smith & Karpinski, 1911; Dauben, 2002). At the same time, integers serve as the foundation for other mathematical representations.
  • Fraction: We consider fractions with integer numerators and denominators. Once the concept of “division” is introduced within the set of integers, the notion of a fraction naturally arises. Extending from integers to fractions (i.e., rational numbers) represents the first major expansion in our understanding of numerical systems. In practical situations, especially those involving distribution, fractions become unavoidable. However, while the concept of fractions is intuitive, performing operations with them is more intricate. As discussed later, mastering fraction operations remains a significant challenge for current LLMs.
  • Floating-Point Number (Float): Floats, or finite decimals, are a useful subset of fractions. Calculations with floats like addition and comparison, work similarly to integers, making them common in daily life. Most fractions are eventually converted to floats when exact precision isn’t needed. Floats are also one of the most prevalent formats in computer. Converting fractions to floats may introduce unacceptable inaccuracies in some cases. Consider the following problem:
    The bookstore sells newspapers and magazines. Bob bought 7 newspapers and 3 magazines for 13 dollars, and Alice bought 3 newspapers and 7 magazines for 23 dollars. How much are the newspapers and magazines, respectively?
    ChatGPT sets up the equations: $137x+3y=$ 13 and $173x+7y=$ 17 and uses the elimination method, $x=\frac{13}{7}-\frac{3}{7}y$ to find the integer solution $x=1$, $y=2$. In this case, the fractions $\frac{13}{7}$ and $\frac{3}{7}$ cannot be approximated as floats (1.857 and 0.429), as this would prevent us from arriving at the exact integer solution and get $y=2.001$ instead. Thus, fractions and floats are tested separately.
  • Scientific Notation: Scientific notation is a common way to represent floats, characterized by separating a number’s precise value from its order of magnitude. The magnitude is expressed using a base-10 exponent, while the float (called significand) represents the precise value. It is widely used in fields like physics, economics, and computer science because it efficiently handles a wide range of numbers and clearly conveys significant figures and precision. For LLMs, mastering scientific notation can significantly enhance their ability to handle practical tasks, such as interpreting financial reports or understanding scientific texts. However, this format introduces complexity, as numbers with different exponents cannot be easily aligned by the decimal point.
@@ 2.2 Tasks in four ability categories
Another aspect of NUPA is defining the tasks that the model needs to handle. The tasks should have clear calculation rules — any student who has completed a full primary and secondary education should be able to accomplish them. Furthermore, most practical numerical processing tasks should either fall within these tasks or can be easily transformed into some of them. We propose 17 tasks across four ability categories. The complete task list is shown in Table 1 and we provide an example for each task in Appendix A.1.2.
Below we discuss the rationales for including some tasks in detail.

### v2

@@ Abstract
Large language models (LLMs) can solve an increasing number of complex reasoning tasks while making surprising mistakes in basic numerical understanding and processing (such as $9.11>9.9$). The latter ability is essential for tackling complex arithmetic and mathematical problems and serves as a foundation for most reasoning tasks, but previous work paid little attention to it or only discussed several restricted tasks (like integer addition). In this paper, we comprehensively investigate the numerical understanding and processing ability (NUPA) of LLMs. Firstly, we introduce a benchmark covering four common numerical representations and 17 distinct numerical tasks in four major categories, resulting in 41 meaningful combinations in total. These tasks are derived from primary and secondary education curricula, encompassing nearly all everyday numerical understanding and processing scenarios, and the rules of these tasks are very simple and clear. Through the benchmark, we find that current LLMs fail frequently in many of the tasks. To study the problem, we train small models with existing and potential techniques for enhancing NUPA (such as tokenizers, PEs, and number formats), comprehensively evaluating their effectiveness using our testbed. We also finetune practical-scale LLMs on our proposed NUPA tasks and find that 1) naive finetuning can improve NUPA a lot on many but not all tasks, and 2) surprisingly, techniques designed to enhance NUPA prove ineffective for finetuning pretrained models. We further explore the impact of chain-of-thought techniques on NUPA. Our work provides a more detailed and comprehensive understanding of NUPA in LLMs. Our benchmark and code are released at https://github.com/GraphPKU/number_cookbook.
@@ 1 Introduction
The mathematical and reasoning abilities of large language models (LLMs) are currently quite impressive (OpenAI, 2023; Meta, 2024a; OpenAI, 2024a; Yang et al., 2024a), capable of solving problems at the level of a graduate student or even more difficult ones like olympiad-level problems (He et al., 2024), GAOKAO (a nationwide examination of high school students applying to universities in China) (Zhang et al., 2024b) and college mathematics (Tang et al., 2024). However, upon closer examination of the models’ outputs, we found that although the models demonstrate remarkable proficiency in problem-solving approaches, they often struggle with basic numerical understanding and processing — like a careless student who claims, “I know how to do it, but I didn’t get it right.” Some of these errors are quite surprising, such as believing that $9.11>9.9$ or making mistakes in simple addition $8/7+3/5$. These errors are a major cause of hallucinations when dealing with math, reasoning, and data analysis tasks, as the model presents seemingly correct problem-solving approaches, but ultimately produces incorrect results (Huang et al., 2024; Li et al., 2024b; Jiang et al., 2024b). Therefore, investigating and improving the fundamental “numerical understanding and processing abilities” (NUPA) of models is crucial.
@@
A more severe issue is that the numerical content is often deliberately simplified in these datasets. In various exam questions (like in the American Invitational Mathematics Examination (Li et al., 2024a)), to focus on assessing students’ understanding of mathematical concepts — such as how to set up the correct equations and apply the right theorems — the numbers in both the questions and answers are often specially chosen to be integers. However, this is not the case in real-world scenarios (Chen et al., 2022a).
@@
On this benchmark, we rigorously test several state-of-the-art LLMs containing GPT-4o (OpenAI, 2024a), Llama-3.1 (Meta, 2024a) and Qwen2 (Qwen Team, 2024). We ask models to directly output the answers without calling external tools. Although the latest LLMs perform well on some easiest tasks, their performance declines significantly as tasks become slightly more complex (such as multiplication, modulo operations, or digit-based calculations), or as the representation of numbers extends beyond basic integers. See Figure 2 of Section 2.4. The overall unsatisfactory performance highlights a pronounced mismatch between the claimed strong mathematical reasoning abilities and the poor practical, everyday numerical understanding and processing abilities of today’s LLMs.
@@
Finally, we discuss the potential of using chain-of-thought (CoT) techniques (Wei et al., 2023) for numerical processing. Although CoT methods can break down complex problems into simpler sub-tasks and significantly increase the likelihood of obtaining correct answers, their drawbacks — such as consuming a large context window and requiring extended processing time — become particularly apparent in numerical tasks. We test a general CoT method known as RFFT (Hu et al., 2024), and find that for more complex tasks (such as multiplication and fraction addition), chain-of-thought methods face scalability challenges, making them difficult to be applied in practical scenarios. It is noteworthy that in this paper, we do not discuss tool use methods (Schick et al., 2023; Lu et al., 2023a) for NUPA as 1) we want to study the self-contained NUPA of LLMs, 2) calling external tools whenever encountering numbers increases the inference latency (Xu et al., 2024), and 3) we believe NUPA without tools is a necessary ability of AGI.
@@ 2 NUPA test: a benchmark for number understanding and processing ability
In this section, we will introduce our NUPA benchmark from the following four aspects: number representations, tasks, metrics, and result analysis of current LLMs. We will explain the rationale behind the inclusion (or exclusion) of specific representations and tasks in our benchmark, highlighting their distinctive features.
@@ 2.1 Number representation
As we discuss above, we believe that the educational curricula on the (Chinese) primary and secondary school levels is a valuable reference for determining the essential NUPAs that LLMs should master. We identify four number formats in these curricula that are both common and sufficient to cover most practical scenarios.
  • Integer: The most common number and the foundation of other number representations.
  • Floating-Point Number (Float): Floats, or finite decimals, are a useful subset of fractions. Calculations with floats like addition and comparison, work similarly to integers, making them common in daily life.
  • Fraction: We consider fractions with integer numerators and denominators. In practical situations involving distribution, fractions become unavoidable, especially when the inaccuracy introduced by converting fractions to floats is unacceptable.
  • Scientific Notation: Scientific notation is characterized by separating a number’s precise value from its order of magnitude. It is widely used in fields like physics, economics, and computer science because it efficiently handles a wide range of numbers and clearly conveys significant figures and precision. For LLMs, mastering scientific notation can significantly enhance their ability to handle practical tasks, such as interpreting financial reports or reading scientific texts.
@@ 2.2 Tasks in four ability categories
Another aspect of NUPA is defining the tasks that the model needs to handle. The tasks should have clear calculation rules — Because these tasks are extracted from the education curricula, students who have completed the stage of education are expected to solve them. Furthermore, most practical numerical processing tasks should either fall within these tasks or can be easily transformed into some of them. We propose 17 tasks across four ability categories. The complete task list is shown in Table 1 and we provide a more detailed discussion in Appendix A.1.2 an example for each task in Appendix A.1.3. Below we discuss the rationales for including some tasks in detail.
  • Elementary arithmetic: addition, subtraction, multiplication, and division. The most fundamental mathematical operations. For division, we consider three types of related operators: True division, floor division and modulus.
  • Comparison: max and min. Understanding numbers on the concept of “order”.
  • Digit understanding: When we care about a language model’s understanding, processing (and generation) of numbers, digit is a crucial concept, as numbers are not read and processed by the language model as a whole, but rather as a sequence of digits. We specially designed some digit-related tasks to test whether LLMs truly handle digits, including:
    • Count: Count the times that a particular digit occurs in an integer.
  • Conversion between representations: Converting a number to two representations: to float and to scientific notation, as they are frequently used to present final results. These two tasks test whether models can understand the relationship between various numerical formats. In particular, since many tasks present answers as approximate values, we designed a “significant digit” (sig. fig.) task to evaluate a model’s ability to round long numbers to fixed-length significant digits.
2.3 Metrics about NUPA
Measuring the performance of NUPA benchmarks on these tasks is not trivial. “Exact match” accuracy is the golden standard of the performance where the answer is considered as correct when it exactly matches the groundtruth. However, a smoother and more detailed metric is useful to understand the behavior and capabilities of a model. Therefore, we also report the “digit match” and “dlength” (difference of length) metrics, as metrics of digit accuracy and length accuracy respectively. We first split numbers into parts (e.g., integer and decimal parts of a float, numerator and denominator of a fraction) and align the generated answer with the groundtruth digit by digit. Integer parts are aligned from the least significant digit; and the decimal parts of float are aligned from the most significant digit. For “digit match”, we measure the correctness of each digit, with missing digits considered as errors, and report the overall accuracy. For “dlength”, we report the sum of absolute difference in length between each part of the prediction and the groundtruth. Figure LABEL:fig:metrics illustrates these three metrics.
2.4 Performance of current LLMs
We test some commonly used LLMs on our benchmark, including three Llama models: Llama-2-7b, Llama-3.1-8b and Llama-3.1-70b (Meta, 2024a), one of the most popular open-source model families from Meta; Mixtral-8$\times$7B (Jiang et al., 2024a), a strong MoE model; and Qwen2-2B and Qwen-72B (Qwen Team, 2024) which are also open-source models that are believed to have strong math abilities. Finally, we also test state-of-the-art commercial models GPT-4o-2024-08-06 and GPT-4o-mini-2024-07-18 (OpenAI, 2024a). We use prompts to control models to directly output result numbers without relying on external tools or CoT. The prompts used for each model and task are included in Appendix A.2. We select the results of some typical tasks in each category in Figure 2, while the complete results and discussion on all metrics are shown in Appendix A.3. And we also provide an interactive performance page where readers can choose tasks they are interested in. Here, we mainly focus on the zero-shot performance while we discuss about few-shot performance in Appendix A.3.1. We have several observations regarding the results:
Figure 2: Parts of performance of state-of-the-art LLMs on NUPA benchmark. “-ft” denotes a Llama model we finetuned on these tasks. (See Section 3.4)
“Digit” is more challenging than expected. We were surprised to find that LLMs struggle to fully grasp “digits”. For instance, in the “get digit” task, where the model is asked to return the $i$-th digit of a long integer, performance drops significantly as the length of the number increases. This suggests that current LLMs lack a consistent ability to just find a digit. Note that the performance is good in the shorter S-range, which indicates that the models can at least comprehend the task instruction. In the XL-range, GPT-4o achieves only 20% accuracy, barely above the random guessing 10% baseline (since the correct answer is always a digit between 0 and 9). This fundamental limitation may explain why current LLMs struggle with numerical understanding and processing, especially as task complexity and input length increase. If a model cannot reliably identify a specific digit in a given number, it casts doubt on its ability to generalize to more complex arithmetic tasks, such as addition.
We also have some interesting observations: (1) LLMs find the “max-hard” task easier than “max” with integer inputs. The difference between these tasks is that in the max task, the two numbers often differ in length, whereas in max-hard, they are always the same length and share some left-most digits, requiring more digits to be compared. While max-hard intuitively seems more difficult, models actually perform better on it. This is likely because they struggle to effectively use sequence length information, as reflected in their weaker performance on the “length” tasks in the longer ranges. It suggests that models might process tasks in different ways from humans. They could have to compare two numbers digit by digit. In this situation, the “harder” subtasks are actually easier because the numbers are already aligned. (2) GPT-4o and GPT-4o-mini show nearly identical performance across most tasks, similar to the comparison between Qwen2-72B and Qwen2-7B. This suggests that once a model reaches a certain size, NUPA performance relies more on factors like architecture design, training strategies, data diversity, and post-training refinements, rather than simply on increasing model size.
3 How do tokenizers, PEs and data formats affect NUPA?
We have observed that the NUPA test poses significant challenges even for the most advanced LLMs. In this section, we aim to investigate the factors that can influence the NUPA of LLMs during their pretraining phase including tokenization strategies, PEs, and different data formats. We utilize the architecture of decoder-only transformers and alter the size to create models with 0.1B, 0.9B and 3B parameters. These models are trained from scratch, incorporating a wide range of techniques that could potentially impact NUPA. In this section, each model is trained on a single task . The details of the training process and models are included in Appendix A.4.1.
3.1 Tokenizer: One-digit tokenizers are good enough
There is a growing tendency to expand the vocabulary size as the number of parameters in LLMs rapidly increases. Recent work has shown that a larger vocabulary is more suitable for larger LLMs (Tao et al., 2024) because longer tokens can encapsulate more complex and precise meanings for text tokens. But number is a different question:
  • The long-tail phenomenon (Raunak et al., 2020), common in text tokens, is not as pronounced for the number tokens. The distribution of number tokens is closer to a uniform distribution.
  • Two smaller number tokens can always be combined into a valid new one (e.g., 3 and 7 form 37), which is not true for text tokens (e.g., “hello” and “hi” cannot form “hellohi”). So the number of possible number tokens grows exponentially as $k$ increases, much faster than text tokens.
  • The next token prediction of number tokens is harder than predicting the next text token because number prediction often involves calculation and operations, whereas word mapping tends to be more intuitive.
(a) 0.9B model int add
(b) 0.9B model float add
(c) 0.9B model int multiply
[…]
From the figure, the one-digit tokenizer shows best in-domain performance in these three tasks, while three-digit tokenizer exhibits poor performance. In out-of-domain tests, one-digit tokenizer even exceeds the others by large margins. Tokenizers with an increasing number of digits significantly hinder sub-billion models’ NUPA. We also performed experiments on models of 3 different sizes including 0.1B, 0.9B, and 3B in Appendix A.4.2 and got similar results. Even as the model size increases, the performance of 2- or 3- digit tokenizer improves but remains either similar or worse than that of the one-digit tokenizers. For these experiments, we also report the digit match and dlength results in Appendix A.4.2 Figure 10 and 11, where one-digit tokenizer performs better both on digit learning (larger digit match) and length learning (less dlength). On the contrary, larger vocabularies significantly increase the model size requirements. In conclusion, we find no evidence to support the idea that increasing the vocabulary size improves NUPA performance.
Recently, Sathe et al. (2024) found that the “random tokenizer” (Kudo, 2018; Provilkov et al., 2020) which splits words like "Hello world" into variable tokens such as "He/llo/ world" or "Hell/o/ world" enhances reasoning by introducing variability in generation path. We also test it in number domain and find the random tokenizers consistently outperform their standard counterparts in length generalization, but still fall short of the performance achieved by the one-digit tokenizer. See the details in Appendix A.4.2.
3.2 Special PEs are length regularization
Previous work has suggested that PE could be the key factor (Zhou et al., 2024) of length generalization. To further investigate whether the influence is specific on a certain task, we train 100M models with different PEs: RoPE (Su et al., 2023), NoPE (Kazemnejad et al., 2023a) and Alibi (Press et al., 2022) on four tasks: integer addition, float addition, fraction multiplication (easy) and scientific notation addition respectively. Models are trained on 1-8 lengths (S and M range), then test them on full range (S to XL, 1-20). RoPE, widely used in Llama and its derivatives, is the most classic relative PE. Then alibi, another relative PE, is proposed to address RoPE’s length overfitting issues. NoPE (transformers without PE, relying solely on the causal mask to encode the position information) offers a surprisingly easy way to achieve length generalization. Therefore, we compare these three typical PEs to evaluate the performance on NUPA.
Our results, presented in Figure 13 in Appendix A.4.3, align with conclusions from previous works. Alibi and NoPE demonstrate superior length generalization across various representations and tasks, indicating that the influence of PEs is relatively consistent across these common representations, tasks within the number domain.
Moreover, we aim to characterize further the mechanism underlying these differences. Specifically, we found that RoPE leads the model to learn a length-related shortcut, while Alibi and NoPE act as a form of regularization by avoiding this, thereby preventing length overfitting. For more details, please refer to the appendix A.4.3.
3.3 Data formats help digit alignment
A series of works have proposed specific data formats including reverse formatting, zero padding and index hints. Reverse formatting (Lee et al., 2023; Shen et al., 2023) presents numbers in reverse order from the least significant digit to the most significant one to align with the models’ autoregressive mechanism, simplifying the learning process for addition. Zero padding (Lee et al., 2023; Shen et al., 2023; Zhou et al., 2024; Cho et al., 2024) adds leading zeros to numbers to standardize the lengths of operands, helping models align operands. Index Hints (Zhou et al., 2023) explicitly include positional information in input and output sequences by representing each digit along with its position index prefixing the digit.
While previous work mainly focuses on integer addition or multiplication, we extend the techniques to various tasks in NUPA Test of different number domains. To compare the effects of reverse formatting and zero padding, we demonstrate in Table 16 how the combination of reverse formatting and zero padding impacts length generalization. Reverse formatting, zero padding, and their combination all show better performance than vanilla formats in integer and float addition, while their performance is comparable to each other, suggesting that their functionality largely overlaps. Zero padding provides the function of helping alignment, and reverse formatting also plays a large role in alignment. The previously believed “helping calculation” function of reverse formatting is minor. As for index hint, we find it doesn’t work for our models. We discuss the details of these experiment results and the reasons in Appendix A.4.4.
3.4 Does finetuning improve NUPA performance of LLMs?
The existing techniques aimed at enhancing NUPA have rarely been applied to practical LLMs, mostly being tested on toy models and isolated tasks. This raises the question of whether it is possible to enhance the NUPA capabilities of large models through post-training finetuning. To explore this, we generate training sets ($10^{5}$ samples for each digit and each task) and validation sets for our NUPA tasks, ensuring no overlap with the original test set. We then used them to perform finetuning on a pre-trained model. Specifically, we finetune a Meta-Llama-3.1-8B model with lora (Hu et al., 2022) (rank 128, $\alpha$=32) on a mixed training set comprising all of our NUPA tasks. Remarkably, we find only 800 steps training (about 50M training samples, $\ll 1$ epoch) leads to significant improvement, as shown in Figure 2 with the finetuned model labeled as “Llama-8B-ft”. Though Llama-3.1-8B is not a strong baseline, this finetuned version achieves much better performance. For example, in max, max-hard, add-float and turediv tasks, this model even surpassed or matched GPT-4o, confirming our hypothesis: for many NUPA tasks, the model’s base capacity may not be the main limiting factor, but rather the lack of numerical diversity and task variety in the training data.
4 Is CoT suitable and valid for NUPA?
CoT has been proven to be effective to enhance the capacity of LLMs both theoretically (Feng et al., 2023; Yang et al., 2024b) and experimentally (Wei et al., 2023; OpenAI, 2024b). Thus, we are also interested in whether CoT is the ultimate solution for improving NUPA. Due to the task and representation diversity in our benchmark, it is hard to cover all issues with a single form of CoT. So we adapt a special CoT form called Rule-Following CoT (Hu et al., 2024) (RF-CoT), where LLMs are trained to follow a provided code or pseudo-code that outlines the procedure to solve the task. RF-CoT is capable of handling any problem with a solving procedure that can be broken down into recurrences and basic unit operations, making it well-suited for our benchmark tasks. The detailed introduction with an example of RF-CoT can be found in Appendix A.5.1.
Table 2: Average inference time.
batchsize sec / sample
RF CoT 128 5.625
Direct 128 0.371
Direct 256 0.336
Exact Match Add Float Multiply Fraction Max Scientific Mod Integer
# Digit 5 6 7 2 3 4 38 39 40 6 7 8
RF CoT 1.00$\pm$.00 1.$0$ 0.93$\pm$.01 0.88$\pm$.03 - 1.00$\pm$.00 1.00$\pm$.00 1.00$\pm$.00 0.67$\pm$.05 0.43$\pm$.07 -
GPT-4o 0.78 0.66 0.49 0.53 0.20 0.00 0.37 0.46 0.36 0.01 0.00 0.00
Qwen2-72B 0.62 0.50 0.70 0.05 0.00 0.00 0.96 0.98 0.95 0.03 0.00 0.00
Llama-8B-ft 0.88$\pm$.02 0.79$\pm$.04 0.74$\pm$.04 0.50$\pm$.02 0.20$\pm$.03 0.01$\pm$.00 0.98$\pm$.01 0.97$\pm$.01 0.98$\pm$.01 0.08$\pm$.02 0.05$\pm$.04 0.05$\pm$.04
To evaluate the performance of this CoT method, we finetuned the LLaMA 3.1-8B model on a subset of the NUPA tasks with RF-CoT. During both training and testing, we set a context window of 2000 tokens, with any data exceeding this limit being ignored. Table 3 shows the performance on selected tasks. Accuracy and standard error for RF-CoT and finetuned Llama-3.1-8B are averaged over three runs. For GPT-4o and Qwen2, which are not finetuned, we report single-run accuracy without standard error. Within the context length limit, the rule-following finetuned LLaMA 3.1-8B significantly outperformed GPT-4o and Qwen2-72B as well as the one finetuned without RF-CoT in most situations.
However, it requires a significantly longer context window and causes much slower inference speed compared to directly generating the answer. As shown in Table 3, with the 2000-token limit, CoT can only handle fraction addition involving numbers up to three digits. We provide the maximal digit length within the 2k context window limitation for each task in Appendix A.5.2 to show the context window limitation for complex tasks. As for inference time, Table 2 demonstrates the average inference time for generating each sample using “RF-CoT” and “direct answer” during the NUPA test where both experiments are operated on an A800 GPU. In the table, the “direct answer” with batch size $256$ shares a similar CUDA memory to RF-CoT with batch size $128$. The RF-CoT method is approximately 17 times slower than directly generating the answer, causing an unsustainable burden for such a basic operation that is frequently encountered in solving real-world problems, especially considering that number calculations may only account for a small part of a complex, real-world reasoning problem (such as analyzing a financial report).
5 Related work
We have discussed some related work in the corresponding section. This section highlights some other studies related to NUPA in language models.
Numerical understanding in natural language comprehension
Earlier studies explored numerical reasoning within language comprehension contexts. For example, Dua et al. (2019) introduced a reading comprehension dataset requiring discrete reasoning, such as sorting and addition. Similarly, Ravichander et al. (2019) proposed a benchmark for evaluating quantitative understanding in textual entailment. However, these datasets blend numerical reasoning with broader language understanding tasks, making it challenging to isolate numerical processing abilities.
Probing numerical understanding in LMs
Several works have probed numerical comprehension in encoder models. Wallace et al. (2019) trained probing models to to assess numerical understanding embedded in model representations, while Johnson et al. (2020) extend this conclusion to multi-language settings. Naik et al. (2019) used contrastive tests to evaluate models’ understanding of number magnitudes. Geva et al. (2020) demonstrated that fine-tuning on numerical reasoning data enhances the understanding. Unlike these studies, which focus on embeddings, our work emphasizes generating correct answers in autoregressive models. Recent efforts on such models include Razeghi et al. (2022), who studied few-shot learning correlations between term frequency and performance, and Zhang et al. (2024a), who identified key components in LLMs for basic arithmetic tasks. These works focus on some most classic tasks and our benchmark expands on these by incorporating diverse numerical representations, tasks, and digit ranges, offering a more comprehensive analysis.
Numerical dataset in specific domains
Datasets like those proposed by Spithourakis & Riedel (2018) and Lin et al. (2020) test numerical commonsense reasoning, while others focus on specific contexts, such as financial reasoning (Chen et al., 2022a, b) or tabular data (Akhtar et al., 2023). These works highlight numerical reasoning within specific domains rather than general numerical processing tasks. In contrast, our benchmark targets core numerical understanding, emphasizing tasks decoupled from domain-specific constraints.
Mathematical Reasoning Datasets
Despite its close relationship with NUPA, mathematical reasoning is a broader field involving diverse skills such as task comprehension, equation solving, tool usage, and more (Lu et al., 2023b). While correct numerical processing is a critical component of mathematical reasoning, it is not the entirety of it (Stolfo et al., 2023). Datasets like MathQA (Amini et al., 2019), GSM8k (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), and SVAMP (Patel et al., 2021) focus on math word problems requiring multi-step reasoning and problem-solving. Few works isolate numerical processing from mathematical reasoning. Saxton et al. (2019) introduced a dataset for numerical tasks, such as adding floating-point numbers, but lacked task categorization by difficulty or length. Moreover, mixing numerical and algebraic tasks complicated analyses of pure numerical processing. Our benchmark addresses this gap, offering fine-grained categorization and evaluation of numerical understanding tasks.
6 Conclusion
We investigate NUPA of LLMs and introduce a comprehensive benchmark, the NUPA test, to reveal that numerical problems remain challenging for modern LLMs. Our comprehensive test suite, which includes a variety of numerical representations and tasks, has exposed the surprising vulnerability of LLMs in this fundamental area. To explore ways to improve NUPA, we extend and evaluate previous pre-training techniques on the NUPA benchmark. While direct finetuning on the NUPA tasks does improve the performance, using those tricks specifically designed for NUPA in the finetuning tends to harm it, suggesting that these methods are not easily transferable to practical LLMs. We also explore the potential of chain-of-thought techniques to enhance NUPA and discuss their limitations.
7 Limitation
As a benchmark that specifically focuses on number understanding and processing abilities, we acknowledge that the range of tasks could still be incomplete and biased toward certain aspects. We will continue updating our benchmark, including but not limited to adding new tasks and refining existing ones to ensure appropriate difficulty. Additionally, the number of models we have tested so far is limited, and we plan to include more promising pre-trained models in future evaluations.
Appendix A Appendix
A.1 NUPA Test
A.1.1 Representations
We present the four representations as follows:
A.1.2 Detailed Introduction and Discussion about tasks
In addition to the brief introduction of the 17 tasks in our benchmark, here we provide a detailed discussion on why these tasks are significant and the specific abilities they aim to evaluate.